A charged capacitor is connected with a resistor after how many time constants does energy of capcitor becomes 1/10th of intial value?

Grade:12th pass

1 Answers

Deepika
3 years ago
V=Q^2/C
Q=√C=Q0/√10........equation(i)
Q=Q0e^-t/c........ equation (ii)
Put equation (i) and (ii) equal to each other.
Q0/√1=Q0e^-t/c
Now Q0 and Q0 will cancel out.
So, 1/√10=1/e^t/c
√10=e^t/c
Multiply by 'ln' on both sides.
ln√10=ln e^t/c
1/2ln10=t/c
11/2×2.303= t/c
t=1.55c
Hope it will help